OpenStudy (anonymous):

find all solutions in interval [0, 2pi) in: sinx - cos x = 0

OpenStudy (anonymous):

At what points on the edge of a circle are the x coordinate and y coordinate equal?

OpenStudy (anonymous):

at pi/4 and 5pi/4 so thats the answer?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

for showing work .. what do I have to do though ?

OpenStudy (anonymous):

Well, you could always draw the unit circle and the line y = x, and say the intersection points are those places where sin(x) = cos(x)

OpenStudy (anonymous):

what about that ( -) sign what does it mean thats what throws me off

OpenStudy (anonymous):

Or you could get fancy and multiply both sides by sin(x) + cos(x), in which case you'd find \[ \sin^2(x) - \cos^2(x) = 0 \rightarrow \sin^2(x) - \left[ 1 - \sin^2(x) \right] = 2\sin^2(x) - 1 = 0\] so \[\sin^2(x) = 1/2\] and the rest follows from there. What minus sign are you talking about?

OpenStudy (anonymous):

sin(x) - cos(x) = 0 and sin(x) = cos(x) mean exactly the same thing.

OpenStudy (anonymous):

yea I just realized wow. haha thank you so much. all the problems were harder and that one so simple just confused me